# Relax Image-Specific Prompt Requirement in SAM: A Single Generic Prompt for Segmenting Camouflaged Objects

### 13 Dec 2023 v2

@@ Introduction
Experimental Results and Analysis
Experimental Results. As shown in Tab. 1, we compare GenSAM with approaches that use different supervision methods, including scribble supervision, point supervision, and our newly proposed generic task prompt supervision. Overall, due to varying levels of supervision signals, scribble supervision outperformed point supervision. However, our GenSAM, despite having only one generic task prompt as the universal supervision for the entire dataset, achieved superior performance compared to point supervision in terms of the $M$, $S_{\alpha}$, and $E_{\phi}$ metrics on CHAMELEON. GenSAM approaches the performance level of scribble supervision. This trend is even more pronounced on the more challenging CAMO and COD10K datasets. Remarkably, on the most challenging COD10K dataset, our method even achieves better results in terms of the $S_{\alpha}$ and $E_{\phi}$ metrics under the less supervised generic task prompt supervision, surpassing both scribble supervision and point supervision approaches. This further demonstrates the superiority of GenSAM. Additionally, our method consistently outperforms SAM, SAM-P, SAM-S and CLIP Surgery+SAM, indicating that the improvements of our method stem from its own merits rather than relying solely on the superior segmentation capabilities of SAM. The qualitative results are shown in Fig. 3.
@@
Component Analysis We further analyzed the impact of components in Tab. 2. When all modules are removed, the model became CLIP Surgery+SAM. We use the general task description “the camouflaged animal” for each sample. The performance is weak across various datasets in this case. This result emphasizes the effectiveness of our complete GenSAM approach. In the third row, we add heatmaps as visual prompts during the iterative process. It is noticed that there is a notable performance improvement. It shows the significance of setting heatmaps as visual prompts, although there is still performance gap compared to GenSAM. In the second and fifth rows, we add consensus heatmaps for foreground and background using the chain-of-thought prompting. The comparison with other experiments emphasizes the importance of chain-of-thought for achieving consensus. The last two rows emphasize the importance of using a chain-of-thought to remove background interference. The unusual results on the CHAMELEON dataset are due to its small size, resulting in greater randomness, while the results from the other two larger and more complex datasets match our expectations. In the forth row, we replace our kkv self-attention with the vvv self-attention from CLIP Surgery. A notable decrease in performance is observed compared to using kkv, indicating the impact of kkv self-attention. As shown in Fig. 3(e), the performance of the model’s test-time adaptation shows a significant boost within the first 1-6 iterations. Then it gradually stabilizes thereafter. Although the best performance is achieved at the 8th iteration, the improvement compared to the 6th iteration is not substantial, and it incurs additional time loss. Therefore, we set the number of iterations to 6.
@@
Generalization of GenSAM. In Tab. 4, we assessed GenSAM’s performance on two segmentation tasks: Polyp Image Segmentation using the ETIS dataset (Silva et al. 2014) and Shadow Detection using the SBU dataset (Vicente et al. 2016). We employed the generic prompts ”Polyp” and ”Shadow” for these respective tasks. Notably, CLIP Surgery + SAM struggled due to the inherent difficulty of distinguishing these objects and the generality of the generic text descriptions. In contrast, our GenSAM boosts the performance greatly. Experiments show the significant improvement achieved by our method compared to the baseline.
Number of chains. We also evaluated the impact of the number of chains in the chain-of-thought prompting in Tab. 3(a). When the number of chains is equal to or less than 3, performance gradually improves with an increasing number of chains. However, when the number of chains exceeded 3, although the inference time increases, there is no significant improvement in performance or even a decline.
Hyper-parameter sensitivity. In Tab.3(b)-3(c) we conduct ablation studies by varying the value of heatmap scale, heatmap threshold under our generic task prompt supervision setting. Four metrics $M$, $F_{\beta}$, $E_{\phi}$ and $S_{\alpha}$ are all adopted to evaluate the robustness. GenSAM is robust to variations in these parameters, validating the stability.
Post processing. We try different methods to transform the mask output result from the last iteration into an auxiliary mask or box prompt in addition to the point prompt to help the current iteration, which ensures the consistency of mask outputs in different iterations. We try different transformation methods, including directly using the last mask output as the mask prompt (Mask), using the maximum surrounding box (MaxBox) and the box that has the largest Intersection over Union of the mask (MaxIOUBox). Tab.3(d) shows that MaxIOUBox outperforms other methods and is the only method that is higher than the baseline in 3 metrics.
@@ Conclusion
In this work, we present GenSAM, which automatically generates image-specific consensus prompts for WSCOD with only a generic task description via our devised test-time progressive mask generation framework. Experiments on diverse COD datasets show the superiority of GenSAM.

### 18 Dec 2023 v3

@@ Introduction
Experiment Results and Analysis
Experiment Results. As shown in Tab. 1, we compare GenSAM with approaches that use different supervision methods, including scribble supervision, point supervision, and our newly proposed generic task prompt supervision. Overall, due to varying levels of supervision signals, scribble supervision outperformed point supervision. However, our GenSAM, despite having only one generic task prompt as the universal supervision for the entire dataset, achieved superior performance compared to point supervision in terms of the $M$, $S_{\alpha}$, and $E_{\phi}$ metrics on CHAMELEON. GenSAM approaches the performance level of scribble supervision. This trend is even more pronounced on the more challenging CAMO and COD10K datasets. Remarkably, on the most challenging COD10K dataset, our method even achieves better results in terms of the $S_{\alpha}$ and $E_{\phi}$ metrics under the less supervised generic task prompt supervision, surpassing both scribble supervision and point supervision approaches. This further demonstrates the superiority of GenSAM. Additionally, our method consistently outperforms SAM, SAM-P, SAM-S and CLIP Surgery+SAM, indicating that the improvements of our method stem from its own merits rather than relying solely on the superior segmentation capabilities of SAM. The qualitative results are shown in Fig. 3.
@@
Component Analysis We further analyze the impact of components in Tab. 2. When all modules are removed, the model became CLIP Surgery+SAM. We use the general task description “the camouflaged animal” for each sample. The performance is weak across various datasets in this case. This result emphasizes the effectiveness of our complete GenSAM approach. In the third row, we add heatmaps as visual prompts during the iterative process. It is noticed that there is a notable performance improvement. It shows the significance of setting heatmaps as visual prompts, although there is still performance gap compared to GenSAM. In the second and fifth rows, we add consensus heatmaps for foreground and background using the chain-of-thought prompting. The comparison with other experiments emphasizes the importance of chain-of-thought for achieving consensus. The last two rows emphasize the importance of using a chain-of-thought to remove background interference. The unusual results on the CHAMELEON dataset are due to its small size, resulting in greater randomness, while the results from the other two larger and more complex datasets match our expectations. In the forth row, we replace our k-k-v self-attention with the v-v-v self-attention from CLIP Surgery. A notable decrease in performance is observed compared to using k-k-v, indicating the impact of k-k-v self-attention. As shown in Fig. 3(e), the performance of the model’s test-time adaptation shows a significant boost within the first 1-6 iterations. Then it gradually stabilizes thereafter. Although the best performance is achieved at the 8th iteration, the improvement compared to the 6th iteration is not substantial, and it incurs additional time loss. Therefore, we set the number of iterations to 6.
@@
Generalization of GenSAM. In Tab. 4, we assess GenSAM’s performance on two segmentation tasks: Polyp Image Segmentation and Shadow Detection. We employ the generic prompts “Polyp” and “Shadow” for them respectively. Experiments demonstrate the significant improvement achieved by our method compared to the baseline.
Number of chains. We also evaluate the impact of the number of chains in the chain-of-thought prompting in Tab. 3(a). When the number of chains is equal to or less than 3, performance gradually improves with an increasing number of chains. However, when the number of chains exceeded 3, although the inference time increases, there is no significant improvement in performance or even a decline.
From heatmap to point prompt. In Tab. 3(b)-3(c) we performe a parameter scan of heatmap upsample factor, the threshold to get the point prompts from heatmap. As the original heatmap $SI$ is of relatively low resolution (14×14)14$\times$ 14)14 × 14 ), we upsample $SI$ to obtain $H$ and obtain the point prompt using a threshold. We finally set the heatmap upsample factor as 2 and the threshold as 0.9
Generating box prompt. In Tab. 3(d), we try different methods to transform the mask output from the previous iteration into an auxiliary mask or box prompt in addition to the point prompt to guide the current iteration, which ensures the consistency of mask outputs in different iterations. We try different transformation methods, including directly using the last mask output as the mask prompt (Mask), using the maximum surrounding box (MaxBox) and the box that has the largest Intersection over Union of the mask (MaxIOUBox). Results show that MaxIOUBox outperforms other strategies
@@ Conclusion
In this paper, we present GenSAM, which automatically generates image-specific consensus prompts for WSCOD with only a generic task description via our test-time progressive mask generation framework. Experiments on various COD datasets show GenSAM’s superiority.
Acknowledgements. This work was supported by Veritone, the Alan Turing Institute Turing Fellowship, and the China Scholarship Council.